# Hermitian and non-Hermitian topology in active matter

### v1

@@ 3.3.5 Non-Bloch band theory
Figure 12: Energy eigenvalues of the non-Hermitian Su-Schrieffer-Heeger model. We show the energy bands obtained from the conventional Bloch band theory (blue curve), the continuum energy bands calculated from Eq. (60) (red curve), and the discrete energy eigenvalues of the finite-size system with the open boundary condition (black dots). The green dots around $E=0$ represent the energies of the topological edge states. The system parameters are set to be $t_{1}=0.5,t_{2}=1$, and $\gamma=2/3$.
As shown in Fig. 10(a), the energy eigenvalues of the open-boundary system can be considerably different from the energy band obtained by the conventional Bloch Hamiltonian (46). Indeed, Eq. (46) reproduces the energy eigenvalues under periodic boundary conditions. This fact indicates that if one invokes the usual Bloch band theory, the extreme boundary sensitivity can induce the breakdown of the conventional bulk-edge correspondence in non-Hermitian systems [204, 205, 206, 207, 208]. Thus, it is necessary to accurately calculate the energy bands that reproduce the energy eigenvalues under open boundary conditions to investigate the topological nature of open-boundary non-Hermitian systems.
Motivated by such observation, the novel type of the Bloch band theory called the non-Bloch band theory has been established as a powerful tool to study non-Hermitian lattice systems [160, 209, 210, 211]. The non-Bloch band theory allows us to calculate the generalized Brillouin zone $\beta=e^{ik}$ for a complex-valued Bloch wavenumber $k$, which gives the continuum sets of the energy eigenvalues under open boundary conditions. The complex-valued Bloch wavenumber indicates that all the eigenstates are localized at the edges due to the non-Hermitian skin effect. Meanwhile, in the conventional Bloch band theory, the Bloch wavenumber takes real values, which means that the bulk eigenstates extend over a whole system. To be concrete, we illustrate the main idea of the non-Bloch band theory by considering the non-Hermitian SSH model. This model is a non-Hermitian extension of the SSH model by introducing intracell asymmetric hopping amplitudes, and its real-space Hamiltonian reads
@@
where $t_{1,2}$ and $\gamma$ take real values. From the general theory of difference equations, one can take an ansatz of the real-space Schrödinger equation, ${\cal H}_{\rm nHSSH}\left|\psi\right\rangle=E\left|\psi\right\rangle$. Specifically, the $x$-site component of the wavefunction can be written as a linear combination of the solutions of the eigenvalue equation, $\det\left[H_{\rm nHSSH}\left(\beta\right)-E\right]=0$, where $H_{\rm nHSSH}\left(\beta\right)$ is given by
@@
Conventionally, the energy eigenvalues of the finite-size system are calculated by combining the eigenvalue equation and the open boundary condition that the wavefunction satisfies. Meanwhile, the non-Bloch band theory allows us to readily calculate the asymptotic continuum sets of the energy eigenvalues in the thermodynamic limit, avoiding the cumbersome calculation process. To explain the method, let $\beta_{1}$ and $\beta_{2}$ denote the solutions of the eigenvalue equation, a quadratic equation of $\beta$. Notably, one can calculate the continuum energy bands by combining the eigenvalue equation and the continuum set of $\beta_{1}$ and $\beta_{2}$, satisfying
@@
The closed curve determined by Eq. (60) is called the generalized Brillouin zone, and it forms the circle with the radius $\sqrt{\left|\left(t_{1}-\gamma\right)/\left(t_{1}+\gamma\right)\right|}$ enclosing the origin of the complex plane. We note that the generalized Brillouin zone becomes a unit circle in Hermitian cases because the Bloch wavenumber takes real values. We show the continuum energy bands calculated from Eq. (60), the energy bands obtained from the conventional Bloch band theory, and the energy eigenvalues of the finite-size system with the open boundary condition in Fig. 12. One can confirm that the continuum energy bands (red curve) truly reproduce the energy eigenvalues (black dots) except for the energies of the topological edge states around $E=0$ (green dots). We remark that the continuum energy band distributes inside the energy band calculated from the real-valued Bloch wavenumber (blue curve).
The continuum energy bands of the non-Hermitian SSH model open the line gap, whose topology can be characterized by the $Z$ topological invariant, similar to the conventional SSH model. Meanwhile, the point-gap topology for the continuum energy bands becomes trivial because the energy winding number takes zero. We next review how to investigate the line-gap topology from the viewpoint of the non-Bloch band theory. According to Refs. [160, 209], the winding number can be defined from the generalized Brillouin zone as follows:
@@
where $\left[{\rm arg}R_{\pm}\left(\beta\right)\right]_{C_{\beta}}$ means the change of ${\rm arg}R_{\pm}\left(\beta\right)$ as $\beta$ goes along the closed curve formed by $\beta=e^{ik}$ in a counterclockwise way. As shown in Figs. 13(a) and (b), the nonzero winding number fully corresponds to the existence of the topological edge states around $E=0$. Meanwhile, one can also define the winding number from the conventional Brillouin zone $e^{ik}~{}\left(k\in R\right)$, similar to Eq. (42). However, as can be seen in Figs. 13(a) and (c), the bulk-edge correspondence appears to be broken in this case, and the winding number calculated from the conventional Brillouin zone cannot be used to predict the appearance of the topological edge states. Thus, the non-Bloch band theory is essential to investigate the line-gap topology of non-Hermitian systems.
@@
Additionally, Ref. [212] has shown that the generalized Brillouin zones are composed of the multiple curves formed by Eq. (62), which become closed curves encircling the origin of the complex plane and can possess some cusps where the curves are indifferentiable. We remark that the existence of the cusps indicates a novel type of topological phase transition. Indeed, Ref. [213] has found that topological semimetal phases appear as intermediate phases between topologically trivial and nontrivial phases, and the exceptional points are stable against perturbations in the topological semimetal phase due to the cusps. Equation (62) can be understood as the condition that the localization lengths of the “plane waves” expressed by $\beta_{M}$ and $\beta_{M+1}$ are equivalent. Said differently, the condition for the generalized Brillouin zone means the formation condition of the standing wave by the plane waves. It is worthnoting that the generalized Brillouin zone and the corresponding continuum energy bands are independent of boundary conditions of open-boundary systems despite the spatial localization of the bulk eigenstates at the edges.
Figure 13: Energy eigenvalues and winding numbers of the non-Hermitian Su-Schrieffer-Heeger model. (a) Energy eigenvalues of the finite-size system with the open boundary condition (black) and energies of the topological edge states (green). (b) Winding number defined in Eq. (61). (c) Winding number calculated from the conventional Brillouin zone $e^{ik}~{}\left(k\in R\right)$. The system parameters are set to be $t_{2}=1$ and $\gamma=2/3$.
@@ 3.3.6 Application of the non-Bloch band theory
which are distinct from the non-Bloch band theory discussed above. The physical meaning of Eq. (63) is that the standing wave formed by $\left|\beta_{2M-1}\right|=\left|\beta_{2M}\right|$ is localized at the edge opposite to one formed by $\left|\beta_{2M}^{-1}\right|=\left|\beta_{2M-1}^{-1}\right|$. Thus, this result is consistent with the localization nature of the symmetry-protected non-Hermitian skin effect.
@@
where $\psi\left(x\right)$ is a wave function, and $p\left(x\right)$, $\lambda_{1,2}\left(x\right)$, and $v\left(x\right)$ are complex-valued periodic functions with a period $a$. The first-order derivative terms in Eq. (64) essentially contribute to realizing the non-Hermitian skin effect since the terms can be regarded as an imaginary gauge potential [177]. The main result of the previous work is that the generalized Brillouin zone becomes the circle with the radius $r$ given by
[…]
which means that the imaginary part of the complex-valued Bloch wavenumber is obtained by $-\frac{1}{a}\log r$. The generalized Brillouin zone allows us to investigate the bulk-edge correspondence, for example, in photonic crystals with anisotropy and optical loss. The electromagnetic wave in the photonic crystal obeys the wave equation written in the form of Eq. (64) and exhibits the non-Hermitian skin effect. The topological edge states appear when the Zak phase defined by the generalized Brillouin zone takes a nonzero value. Additionally, Ref. [216] has investigated the construction of the non-Bloch band theory in spatially periodic continuous systems described by generalized eigenvalue problems.
The relaxation dynamics from initial states to a steady state have been studied in stochastic processes described by Eq. (52). The steady state uniquely corresponds to the eigenvector of the transition matrix whose eigenvalue is equal to 00. We remark that the non-Bloch band theory is useful for investigating the role of topology in the relaxation dynamics. Let us suppose that the system has the spatial translation symmetry, and the transition matrix is denoted by $W_{nm}$, where $n$ and $m$ are the site indices. Furthermore, let $W\left(k\right)$ denote the diagonalized form of the transition matrix in the Fourier basis. For the sake of convenience, we introduce the scale-transformation form of the transition matrix as follows: $\left(W^{\lambda}\right)_{nm}\equiv W_{nm}e^{\lambda\left(n-m\right)}$. Reference [217] has proposed that the point-gap topology can be characterized by the winding number defined as
@@
[…]
The main claim of the previous work is that, when the winding number takes a nonzero value, the spectrum obtained by the non-Bloch band theory has a finite gap below the zero eigenvalue [Fig. 14]. Therefore, this claim indicates that in a finite-size system with open boundary conditions the non-Hermitian skin effect occurs, and an eigenvalue of the transition matrix is either within the continuum spectrum or equal to zero. Importantly, the existence of the gap leads to a finite relaxation time in the stochastic process in an intuitive sense. However, the relaxation time also depends on the overlap between initial states and the eigenvectors of the transition matrix [218, 219]. One can immediately infer that the overlap decreases because the eigenvectors are localized due to the non-Hermitian skin effect. As a result, both the inverse of the gap and of the overlap contribute to the relaxation time, and the relaxation time becomes proportional to the system size. This system-size dependence of the relaxation time is still different from that when the winding number vanishes, because it is proportional to the square of the system size in a topologically trivial system.
Finally, we comment on recent efforts in extending the non-Bloch band theory to 2D non-Hermitian systems. Reference [220] has shown that, in 2D tight-binding systems exhibiting the non-Hermitian skin effect in only one direction, one can determine the generalized Brillouin zone and calculate the continuum energy bands on the square geometry. As a result, the bulk-edge correspondence has been successfully established, for example, in the Chern insulator with the asymmetric hopping amplitude. The previous work has also pointed out that the effect of the shape of the systems on the standing wave leads to the difficulty of calculating the continuum energy bands. Reference [221] has proposed the condition for the generalized Brillouin zone of 2D non-Hermitian systems, which goes beyond Ref. [220]. However, the validity of the calculation method has been under debate. Thus, whether or not the non-Bloch band theory in 2D non-Hermitian systems can be genuinely constructed remains an open problem.
@@ 3.3.7 Exceptional point
which depends on two real parameters $\gamma$ and $a$. $i\gamma$ represents the on-site gain and loss, which leads to the non-Hermiticity of the Hamiltonian. If one multiplies the Hamiltonian by the imaginary unit $i$, $\gamma$ corresponds to the natural frequency and $a$ represents the dissipation between two components. This Hamiltonian has exceptional points at $\gamma=\pm a$, where one obtains only one eigenvector $\psi\propto(i,\pm 1)^{T}$. Figure 15(b) shows the eigenvalues of the Hamiltonian at different $\gamma$. We can see the singular behavior around the exceptional points, which plays important roles both in the topological protection of the exceptional point and its application to enhanced sensitivity as we will discuss below.
@@
This Hamiltonian has exceptional points at $k_{x}=k_{y}=0$. Figure 15(c) shows the complex spectra around the exceptional point, which exhibit a characteristic branchpoint structure. The exceptional point is protected by the nontrivial topology of this branchpoint structure around it, and thus cannot be removed until two or more exceptional points coalesce and pair-annihilate. To clearly show its topological feature, one should consider the eigenvalues on a circular path $C$ around the exceptional point. Moving along the path $C$, a pair of eigenvalues are swapped to each other. This braiding cannot be disentangled until other gapless points cross the path $C$, which is similar to the fact that a Möbius band cannot be deformed to a normal band as long as one cuts the band. Meanwhile, the swapping of eigenvalues is unique to gapless systems because if spectra have a gap, one can distinguish and separate each spectrum. Therefore, exceptional points are protected by the topology of the spectra around them including the braiding of eigenvalues. One can also define the topological invariant of exceptional points as
[…]
@@
As in Hermitian topological insulators, exceptional points can also be protected by symmetries. To see this, we again consider the Hamiltonian (69). When we fix $a$ and assume $\gamma$ as a wavenumber $\gamma=k$, the Hamiltonian is considered as a non-Hermitian Bloch Hamiltonian. This one-dimensional Hamiltonian is symmetric under the combination of the parity inversion (i.e., space inversion) and the time reversal, which is known as the parity-time ($PT$) symmetry [224, 225]. By using a unitary matrix $PT$, one can define the $PT$ symmetry of the Hamiltonian as $PTH(k)(PT)^{-1}=H^{\ast}(k)$. The $PT$ symmetry guarantees that eigenvalues are real or have complex-conjugate counterparts. To continuously modify the complex-conjugate pair of eigenvalues into real ones, the Hamiltonian must be degenerate at the transition point. Such a degenerate point corresponds to an exceptional point and cannot be removed until two exceptional points coalesce, which indicates the topological protection of exceptional points under the $PT$ symmetry. In fact, one can define the topological invariant as
@@
As in the gapped topological phases, a topological classification of exceptional points has also been proposed in Ref. [223]. Topological classification of $d$-dimensional exceptional structures (cf. $d=0$ for the exceptional point and $d=1$ for the exceptional ring) in $D$-dimensional systems corresponds to the topological classification of gapped phases in $D-d-1$ dimension. However, we need to consider $PT$ and $CP$ symmetries instead of the conventional time-reversal and charge-conjugate symmetries, as we have already seen in the one-dimensional case.
@@
where the diagonal terms ±$\pm k$ represent the linear dispersion of edge modes in each layer of a Chern insulator and the off-diagonal terms $-ia$ correspond to the non-Hermitian coupling. Since this effective Hamiltonian is equal to the Hamiltonian (69) multiplied by $-i$, the Hamiltonian exhibits gapless dispersions with exceptional points. One can also confirm the topological protection of gapless bands in the same way as exceptional points in one-dimensional bulk bands. The exceptional boundary modes can be realized in photonic systems [246], and one can also find them in active matter [245] as discussed below.

### v2

@@ 4.3.3 Non-Bloch band theory
Figure 12: Energy eigenvalues of the non-Hermitian Su-Schrieffer-Heeger model. We show the energy bands obtained from the conventional Bloch band theory (blue curve), the continuum energy band calculated from Eq. (61) (red curve), and the discrete energy eigenvalue of the finite-size system with an open boundary condition (black dots). The green dots around $E=0$ represent the energies of the topological edge states. The system parameters are set to be $t_{1}=0.5,t_{2}=1$, and $\gamma=2/3$.
As shown in Fig. 10(a), the energy eigenvalues of the open-boundary system can be considerably different from the energy band obtained by the conventional Bloch Hamiltonian (47). Indeed, Eq. (47) reproduces the energy eigenvalues under periodic boundary conditions. Importantly, if one invokes the usual Bloch band theory, the extreme boundary sensitivity can induce the breakdown of the bulk-edge correspondence associated with the line-gap topology [212, 213, 214, 215, 216]. To investigate the topological nature of open-boundary non-Hermitian systems, it is necessary to accurately calculate the energy band that reproduces the energy eigenvalue under open boundary conditions.
Motivated by such observation, the novel type of the Bloch band theory called a non-Bloch band theory has been established as a powerful tool to study non-Hermitian lattice systems [165, 217, 218, 219]. The non-Bloch band theory allows us to calculate the generalized Brillouin zone $\beta=e^{ik}$ for a complex-valued Bloch wavenumber $k$, which gives the continuum sets of the energy eigenvalues under open boundary conditions. The complex-valued Bloch wavenumber indicates that all the eigenstates are localized at the edge due to the non-Hermitian skin effect. Meanwhile, in the conventional Bloch band theory, the Bloch wavenumber takes real values, which means that all the bulk eigenstates extend over a whole system. To be concrete, we illustrate the main idea of the non-Bloch band theory by considering a non-Hermitian SSH model. This model is a non-Hermitian extension of the SSH model (cf. Fig. 8) by introducing intracell asymmetric hopping amplitudes, and its real-space Hamiltonian reads
@@
where $t_{1,2}$ and $\gamma$ take real values. From a general theory of difference equations, one can take an ansatz of the real-space Schrödinger equation, ${\cal H}_{\rm nHSSH}\left|\psi\right\rangle=E\left|\psi\right\rangle$. Specifically, the $x$-site component of the wavefunction can be written as a linear combination of the solutions of the eigenvalue equation, $\det\left[H_{\rm nHSSH}\left(\beta\right)-E\right]=0$, where $H_{\rm nHSSH}\left(\beta\right)$ is given by
@@
Conventionally, the energy eigenvalue of the finite-size system is calculated by combining the eigenvalue equation and open boundary conditions that the wavefunction satisfies. Meanwhile, the non-Bloch band theory allows us to readily calculate the asymptotic continuum sets of the energy eigenvalues in the thermodynamic limit, avoiding the cumbersome calculation process. To explain the method, let $\beta_{1}$ and $\beta_{2}$ denote the solutions of the eigenvalue equation, a quadratic equation of $\beta$. Notably, one can calculate the continuum energy band by combining the eigenvalue equation and the continuum set of $\beta_{1}$ and $\beta_{2}$, satisfying
@@
The closed curve determined by Eq. (61) is called the generalized Brillouin zone, and it forms the circle with the radius $\sqrt{\left|\left(t_{1}-\gamma\right)/\left(t_{1}+\gamma\right)\right|}$ enclosing the origin of the $\beta(=e^{ik})$ plane. We note that the generalized Brillouin zone becomes a unit circle in Hermitian cases because the Bloch wavenumber takes real values. We show the continuum energy bands calculated from Eq. (61), the energy bands obtained from the conventional Bloch band theory, and the energy eigenvalues of the finite-size system with an open boundary condition in Fig. 12. One can confirm that the continuum energy bands (red curve) truly reproduce the energy eigenvalues (black dots) except for the energies of the topological edge states around $E=0$ (green dots). We remark that the continuum energy bands are distributed inside the energy band calculated from the real-valued Bloch wavenumber (blue curve).
The continuum energy bands of the non-Hermitian SSH model open the line gap, whose topology can be characterized by the $Z$ topological invariant, similar to the conventional SSH model. Meanwhile, the point-gap topology for the continuum energy band becomes trivial because the energy winding number always takes zero. We next review how to investigate the line-gap topology from the viewpoint of the non-Bloch band theory. According to Refs. [165, 217], the winding number can be defined from the generalized Brillouin zone as follows:
@@
where $\left[{\rm arg}R_{\pm}\left(\beta\right)\right]_{C_{\beta}}$ means the change of ${\rm arg}R_{\pm}\left(\beta\right)$ as $\beta$ goes along the closed curve formed by $\beta=e^{ik}$ in a counterclockwise way. As shown in Figs. 13(a) and (b), the nonzero winding number fully corresponds to the existence of the topological edge states around $E=0$. Meanwhile, one can also define another winding number from the conventional Brillouin zone $e^{ik}~\left(k\in R\right)$, similar to Eq. (43). However, as can be seen in Figs. 13(a) and (c), the bulk-edge correspondence appears to be broken in this case, and the winding number calculated from the conventional Brillouin zone cannot be used to predict the appearance of the topological edge states. Thus, the non-Bloch band theory is essential to investigate the line-gap topology of non-Hermitian systems.
@@
Additionally, Ref. [220] has shown that the generalized Brillouin zones are composed of the multiple curves formed by Eq. (63), which become closed curves encircling the origin of the $\beta(=e^{ik})$ plane and can possess some cusps where the curves are indifferentiable. We remark that the existence of the cusps indicates a novel type of topological phase transition. Indeed, Ref. [221] has found that a topological semimetal phase appears as an intermediate phase between topologically trivial and nontrivial phases, and gap-closing points (exceptional points) are stable against perturbations in the topological semimetal phase due to the cusps. Equation (63) can be understood as the condition that the localization lengths of “plane waves” expressed by $\beta_{M}$ and $\beta_{M+1}$ are equivalent. Said differently, the condition for the generalized Brillouin zone means the formation condition of the standing wave by the plane waves. It is worth noting that the generalized Brillouin zone and the corresponding continuum energy band are independent of boundary conditions of open-boundary systems despite the spatial localization of the bulk eigenstates at the edge.
Figure 13: Energy eigenvalue and winding numbers of the non-Hermitian Su-Schrieffer-Heeger model. (a) Energies of the bulk modes (black) and energies of the topological edge states (green) in the finite-size system with an open boundary condition. (b) Winding number defined in Eq. (62). (c) Winding number calculated from the conventional Brillouin zone $e^{ik}~\left(k\in R\right)$. The system parameters are set to be $t_{2}=1$ and $\gamma=2/3$.
@@ 4.3.4 Application of the non-Bloch band theory
which are distinct from the condition (63). The physical meaning of Eq. (64) is that the standing wave formed by $\left|\beta_{2M-1}\right|=\left|\beta_{2M}\right|$ is localized at the edge opposite to one formed by $\left|\beta_{2M}^{-1}\right|=\left|\beta_{2M-1}^{-1}\right|$. Thus, this result is consistent with the localization nature of the symmetry-protected non-Hermitian skin effect.
@@
where $\psi\left(x\right)$ is a wave function, and $p\left(x\right)$, $\lambda_{1,2}\left(x\right)$, and $v\left(x\right)$ are complex-valued periodic functions with a period $a$. The first-order derivative terms in Eq. (65) essentially contribute to realizing the non-Hermitian skin effect since the terms can be regarded as an imaginary gauge potential [184]. The main result of the previous work is that the generalized Brillouin zone becomes a circle with a radius
$r=\exp\left(\frac{1}{2}\int_{0}^{a}dx\;{\rm Im}\left[\frac{\lambda_{1}\left(x\right)+\lambda_{2}\left(x\right)}{p\left(x\right)}\right]\right),$ (66)
which means that the imaginary part of the complex-valued Bloch wavenumber is obtained by $\frac{1}{a}\log r$. The generalized Brillouin zone allows us to investigate the bulk-edge correspondence, for example, in photonic crystals with anisotropy and optical loss. The electromagnetic wave in the photonic crystal obeys the wave equation written in the form of Eq. (65) and exhibits the non-Hermitian skin effect. The topological edge state appears when the Zak phase defined by the generalized Brillouin zone takes a nonzero value. Recently, a general extension of the framework has been proposed in Ref. [224]. Additionally, Ref. [225] has investigated the construction of the non-Bloch band theory in spatially periodic continuous systems described by generalized eigenvalue equations.
The relaxation dynamics from initial states to a steady state have been studied in stochastic processes described by Eq. (53). The steady state uniquely corresponds to the eigenvector of the transition matrix whose eigenvalue is equal to 0. We remark that the non-Bloch band theory is useful for investigating the role of topology in the relaxation dynamics. Let us suppose that the system has the spatial translation symmetry, and the transition matrix is denoted by $W_{nm}$, where $n$ and $m$ are the site indices. Furthermore, let $W\left(k\right)$ denote the diagonalized form of the transition matrix in the Fourier basis. For the sake of convenience, we introduce the scale-transformation form of the transition matrix as follows: $\left(W^{\lambda}\right)_{nm}\equiv W_{nm}e^{\lambda\left(n-m\right)}$. Reference [226] has proposed that the point-gap topology can be characterized by the winding number defined as
@@
$w_{\pm}=\lim_{\lambda\rightarrow\pm 0}\frac{1}{2\pi i}\int_{0}^{2\pi}dk\frac{d}{dk}\log\left(\det W^{\lambda}\left(k\right)\right).$ (68)
The main claim of the previous work is that, when the winding number takes a nonzero value, the non-Hermitian skin effect occurs, and the spectrum obtained by the non-Bloch band theory has a finite gap below the zero eigenvalue [Fig. 14]. Therefore, this claim indicates that an eigenvalue of the transition matrix is either within the spectrum or equal to zero in a finite-size system with arbitrary open boundary conditions. Importantly, the existence of the gap leads to a finite relaxation time in the stochastic process in an intuitive sense. However, the relaxation time also depends on the overlap between initial states and the eigenvectors of the transition matrix [227, 228]. One can immediately infer that the overlap decreases because the eigenvectors are localized due to the non-Hermitian skin effect. As a result, both the inverse of the gap and of the overlap contribute to the relaxation time, and the relaxation time becomes proportional to the system size. This system-size dependence of the relaxation time is still different from that when the winding number vanishes, because it is proportional to the square of the system size in a topologically trivial system.
Finally, we comment on recent efforts in extending the non-Bloch band theory to 2D non-Hermitian systems. Reference [229] has shown that, in 2D tight-binding systems with specific symmetries, one can determine the generalized Brillouin zone and calculate the continuum energy band as in the case of 1D tight-binding systems. However, the previous work has pointed out that the effect of systems’ shapes on the formation of the standing wave leads to the difficulty of calculating the continuum energy band. Beyond the previous work, Ref. [230] has proposed an amoeba formulation to calculate the continuum energy band in higher-dimensional tight-binding systems on arbitrary geometry. Importantly, the amoeba formulation allows us to successfully establish the bulk-edge correspondence in the Chern insulator with asymmetric hopping amplitudes. While the previous work has clarified the calculation method for the continuum energy band of the bulk skin mode, the construction of the non-Bloch band theory for higher-order skin mode is still an open problem.
@@ 4.4 Exceptional point
which depends on two real parameters $\gamma$ and $a$. $i\gamma$ represents the on-site gain and loss, which leads to the non-Hermiticity of the Hamiltonian. If one multiplies the Hamiltonian by the imaginary unit $i$, $\gamma$ corresponds to the natural frequency and $a$ represents the dissipation between two components, the latter of which can be introduced by, e.g., aligning interaction between active particles as we will discuss in Sec. 5.3.2. This Hamiltonian has exceptional points at $\gamma=\pm a$, where one obtains only one eigenvector $\psi\propto(i,\pm 1)^{T}$. Figure 15(b) shows the eigenvalues of the Hamiltonian at different $\gamma$. We can see the singular behavior around the exceptional points, which plays important roles both in the topological protection of the exceptional point and its application to enhanced sensitivity as we will discuss below.
@@
This Hamiltonian has exceptional points at $k_{x}=k_{y}=0$. Figure 15(c) shows the complex spectra around the exceptional point, which exhibit a characteristic branchpoint structure. The exceptional point is protected by the nontrivial topology of this branchpoint structure around it, and thus cannot be removed until two or more exceptional points coalesce and pair-annihilate. To clearly show its topological feature, one should consider the eigenvalues on a circular path $C$ around the exceptional point. Moving along the path $C$, a pair of eigenvalues are swapped to each other. This braiding cannot be disentangled until other gapless points cross the path $C$, which is similar to the fact that a Möbius band cannot be deformed to a normal band as long as one does not cut the band. Meanwhile, the swapping of eigenvalues is unique to gapless systems because if spectra have a gap, one can distinguish and separate each spectrum. Therefore, exceptional points are protected by the topology of the spectra around them including the braiding of eigenvalues. One can also define the topological invariant of exceptional points as
$w=\oint_{C}d\mathbf{k}\cdot\nabla_{\bf k}\log\det[H(\mathbf{k})-E_{0}],$ (72)
@@
Due to the topological property of exceptional points characterized by the winding number, exceptional points have robustness against perturbations to the Hamiltonian. On the other hand, when two or more exceptional points coalesce, the gapless path $C$ in Eq. (72) cannot be defined around each exceptional point, and thus the band topology can be altered. In fact, if the sum of the winding numbers is zero, exceptional points can be removed via the coalescence, which resembles the annihilation of topological defects in the real space [13, 14].
As in Hermitian topological insulators, exceptional points can also be protected by symmetries. To see this, we again consider the Hamiltonian (70). When we fix $a$ and assume $\gamma$ as a wavenumber $\gamma=k$, the Hamiltonian is considered as a non-Hermitian Bloch Hamiltonian. This one-dimensional Hamiltonian is symmetric under the combination of the parity inversion (i.e., space inversion) and the time reversal, which is known as the parity-time ($PT$) symmetry [233, 234]. By using a unitary matrix $PT$, one can define the $PT$ symmetry of the Hamiltonian as $PTH(k)(PT)^{-1}=H^{\ast}(k)$. The $PT$ symmetry guarantees that eigenvalues are real or have complex-conjugate counterparts. To continuously modify the complex-conjugate pair of eigenvalues into real ones, the Hamiltonian must be degenerate at the transition point as demonstrated in Fig. 15(a). Such a degenerate point corresponds to an exceptional point and cannot be removed until two exceptional points coalesce, which indicates the topological protection of exceptional points under the $PT$ symmetry. In fact, one can define the topological invariant as
@@
As in the gapped topological phases, a topological classification of exceptional points has also been proposed in Ref. [232]. Topological classification of $d$-dimensional exceptional structures (cf. $d=0$ for the exceptional point and $d=1$ for the exceptional ring) in $D$-dimensional systems corresponds to the topological classification of gapped phases in Dd1D-d-1 dimension. However, we need to consider $PT$ and $CP$ symmetries instead of the conventional time-reversal and charge-conjugate symmetries, as we have already seen in the one-dimensional case.
@@
where the diagonal terms $\pm k$ represent the linear dispersion of edge modes in each layer of a Chern insulator and the off-diagonal terms ia-ia correspond to the non-Hermitian coupling. Since this effective Hamiltonian is equal to the Hamiltonian (70) multiplied by i-i, the Hamiltonian exhibits gapless dispersions with exceptional points. One can also confirm the topological protection of gapless bands in the same way as exceptional points in one-dimensional bulk bands. The exceptional boundary modes can be realized in photonic systems [254], and one can also find them in active matter [124] as discussed later in Sec. 5.3.2.